apple pie is cut into 6 slices. each slice is sold for $2,the radius of the pie is 3.5 inches. central angle of the remaining pie is 120 degree.how much does remaing pie cost
Answer by richwmiller(17219)

120/360=1/3 of the pie is remaining.
6/3=2 slices
2*2=4 dollars